OpenStudy (anonymous):

A consumer initially spends 60% of his income on food. The price of food rises by 20%. At the same time, the individual is compensated by income increase of 12%. Therefore individual must be worse off. True or false,explain

OpenStudy (anonymous):

LET INCOME= 100 & PRICE OF FOOD=10 EXPENDITURE ON FOOD=60 which means UNITS OF FOOD=60/10=6 AFTER PRICE INCREASE N COMPENSATION, PRICE OF FOOD=12 COMPENSATED INCOME= 112 NOW,units of food= 112/12=9.33 THUS, INDVL WIL B BETTER OFF as he is able to get more units of food now